# The second class particle in the half-line open TASEP

### v1

@@ 4.2. Explicit formulas for asymptotic distribution of the second class particles
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
@@
The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(4.12) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.13) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
@@
(4.14) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.15) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
(4.16) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
(4.17) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
@@
The proof directly follows from Theorem 3.2, Theorem 3.3, Theorem 4.2 and Remark 4.8. ∎
@@
The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(4.18) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.19) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
@@
(4.20) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.21) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
(4.22) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
(4.23) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
@@
The proof directly follows from Theorem 3.7, Theorem 4.2 and Remark 4.8. ∎
@@ 5. Asymptotic distribution of second class particles under KPZ-type scaling
provided the series converges, and we recall that for an skew-symmetric $2k\times 2k$ matrix $A$, its Pfaffian is defined by
@@
$\mathbb{P}\left(2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2}\geq 0\right)$

### v2

@@ 4.2. Explicit formulas for asymptotic distribution of the second class particles
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha\leq\frac{1}{2}$:
@@
Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(4.12) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}.$
@@
(4.13) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}.$
@@
The proof directly follows from Theorem 3.3, Theorem 3.4, Theorem 4.1. ∎
Our second main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha>\frac{1}{2}$:
Theorem 4.11 (Modulo Conjecture 4.3).
Let $\alpha>\frac{1}{2}$ and we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(4.14) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
(4.15) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
Proof.
The proof directly follows from Theorem 3.3, Theorem 3.4, Theorem 4.2, Conjecture 4.3 and Theorem 4.4. ∎
Our third main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha\leq\frac{1}{2}$:
@@
Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(4.16) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}.$
@@
(4.17) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}.$
@@
The proof directly follows from Theorem 3.8, theorem 4.1. ∎
Our fourth main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha>\frac{1}{2}$:
Theorem 4.13 (Module Conjecture 4.3).
Let $\alpha>\frac{1}{2}$ and we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(4.18) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
(4.19) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
Proof.
The proof directly follows from Theorem 3.8, Theorem 4.2, Conjecture 4.3 and Theorem 4.4. ∎
@@ 5. Asymptotic distribution of second class particles under KPZ-type scaling
provided the series converges, and we recall that for a skew-symmetric $2k\times 2k$ matrix $A$, its Pfaffian is defined by
@@
$\mathbb{P}\left(2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2}\geq 0\right)$
@@
Assume that $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. The asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\xi_{t}$ is given by: